Jerry’s Window Service received $14,000 from a client on February 20. This payment was an advance payment for 7 months of window cleaning starting March 1. The window cleaning services are provided equally over the 7 months. At May 31, calculate the balance in the unearned revenue account assuming that all adjusting entries have been properly recorded. Select one: $14,000 $ 6,000 $10,000 $ 8,000 On December 1, Optima Corp. paid $6,000 for rent expense covering December, January and February rent. On December 31, after all adjusting entries, Optima Inc. will report Prepaid Rent of: Select one: $4,000 $0 $6,000 $2,000
Answer #1
1
By May 31, 3 months of services have been performed and 4 months are remaining.
Balance in the unearned revenue account = 14000*4/7= $8000
$8000 is the correct option
2
Prepaid Rent will be reported for 2 months
Prepaid Rent = 6000*2/3 = $4000
$4000 is the correct option